# NCL++: Nested Collaborative Learning for Long-Tailed Visual Recognition

### v1

@@ Abstract
Long-tailed visual recognition has received increasing attention in recent years. Due to the extremely imbalanced data distribution in long-tailed learning, the learning process shows great uncertainties. For example, the predictions of different experts on the same image vary remarkably despite the same training settings. To alleviate the uncertainty, we propose a Nested Collaborative Learning (NCL++) which tackles the long-tailed learning problem by a collaborative learning. To be specific, the collaborative learning consists of two folds, namely inter-expert collaborative learning (InterCL) and intra-expert collaborative learning (IntraCL). InterCL learns multiple experts collaboratively and concurrently, aiming to transfer the knowledge among different experts. IntraCL is similar to InterCL, but it aims to conduct the collaborative learning on multiple augmented copies of the same image within the single expert. To achieve the collaborative learning in long-tailed learning, the balanced online distillation is proposed to force the consistent predictions among different experts and augmented copies, which reduces the learning uncertainties. Moreover, in order to improve the meticulous distinguishing ability on the confusing categories, we further propose a Hard Category Mining (HCM), which selects the negative categories with high predicted scores as the hard categories. Then, the collaborative learning is formulated in a nested way, in which the learning is conducted on not just all categories from a full perspective but some hard categories from a partial perspective. Extensive experiments manifest the superiority of our method with outperforming the state-of-the-art whether with using a single model or an ensemble. The code will be publicly released.
@@ 1 Introduction
In recent years, deep neural networks have achieved resounding success in various visual tasks, e.g., image classification he2016deep ; hu2018squeeze , object detection ren2015faster ; zhang2018single , semantic segmentation zhao2017pyramid ; fu2019dual and so on. Despite the advances in deep technologies and computing capability, the huge success also highly depends on large well-designed datasets of having a roughly balanced distribution, such as ImageNet deng2009imagenet , MS COCO lin2014microsoft and Places zhou2017places . This differs notably from real-world datasets, which usually exhibits long-tailed data distributions wang2017learning ; liu2019large where few head classes occupy most of the data while many tail classes only have few samples as shown in Fig. 1. In such scenarios, the model is easily dominated by those few head classes, whereas low accuracy rates are usually achieved for many other tail classes. Undoubtedly, the long-tailed characteristics challenge deep visual recognition, and also immensely hinder the practical use of deep models.
In long-tailed visual recognition, most early works focus on designing the class re-balancing strategies he2009learning ; buda2018systematic ; cui2019class ; huang2016learning ; ren2018learning ; wang2017learning and decoupled learning kang2019decoupling ; cao2019learning . Those methods can accomplish some accuracy improvements but still cannot deal with the long-tailed class imbalance problem well. For example, class re-balancing methods are often confronted with risk of overfitting. More recent efforts aim to improve the long-tailed learning by using multiple experts xiang2020learning ; wang2020long ; li2020overcoming ; cai2021ace ; zhang2021test . The multi-expert algorithms follow a straightforward idea of making them diverse from each other. To achieve this, some works xiang2020learning ; zhang2021test force different experts to focus on different aspects. For example, LFME xiang2020learning formulates a network with three experts and it forces each expert to learn samples from one of head, middle and tail classes. Besides, there are some works wang2020long adopt some constraint losses to diversify the multiple experts. For example, RIDE wang2020long proposes a distribution-aware diversity loss to achieve this. The previous multi-expert methods can achieve good performance mainly due to the diversity among them. Also because of this, these methods obtain the predictions through the ensemble rather than a single expert due to that only employing a single expert hardly achieves reliable recognition.
@@
However, we found that there are great uncertainties in the long-tailed learning. The learning uncertainties mainly come from two aspects. One is that the same expert performs differently for the same image with different augmentations. The other is that, for two experts with the same configuration and the same training settings, they still vary greatly on predictions with respect to the same input image. To analyze this, we visualize the Kullback-Leibler (KL) distance [^1] https://en.wikipedia.org/wiki/Kullback-Leibler_divergence over the predictions generated by two experts with the same training settings, and the predictions generated by two augmented image copies with respect to the same expert. As shown in Fig. 2 (b), the green color shows the KL distance distribution for two different experts of the same structure and training settings with taking the same image as the input, and the blue color shows that of the same expert with taking different augmentations of the same image. Specifically, to analyze the uncertainties across all categories, the Kullback-Leibler (KL) distance is specifically calculated for all categories with reporting the average distance over all images for each class. As we can see, the predictions vary greatly especially in tail classes, which signifies the great uncertainty in the learning process as well as the diversity in models. To show the prediction differences in details, we visualize the detailed predictions of the two experts with the same structure and training settings on a randomly selected sample (corresponding to the green color in Fig. 2 (b)). According to previous works xiang2020learning ; wang2020long ; li2020overcoming ; cai2021ace ; zhang2021test , a simple ensemble like averaging the predictions could be employed to reduce the uncertainty in the prediction stage, which can achieve a certain performance improvement. However, an ensemble of multiple experts also brings huge amount of calculation and parameters, which limits its practical use. This brings up a question, can we utilize these uncertainties to improve the performance of single model?
@@
Previous works guo2020online ; lan2018knowledge ; zhang2018deep simply conduct the classification and distillation from a full perspective on all categories. It helps the network to obtain the global discrimination on all categories, but lacks of meticulous distinguishing ability on the confusing categories. In the classification problem, the network should pay attention to the hard negative categories with high predictive scores rather than simply treating all categories equally, which helps to reduce the confusion between the target category and the confusing categories. To achieve this, we first propose a Hard Category Mining (HCM) to select the hard categories for each sample, in which the hard category is defined as the negative category with a high predictive score. Then, we formulate a nested learning for the supervised individual learning of a single network as well as the collaborative learning (including IntraCL and InterCL). To the specific, the nested learning conducts the supervised learning or distillation from two perspectives, where one is the full perspective on all categories and the other is the partial perspective only on the selected hard categories. In this way, not only the global discriminative capability but also the meticulous distinguishing ability can be captured.
The proposed method utilizes the collaborative learning for long-tailed visual recognition. The proposed collaborative learning is two folds, where one is inter-expert collaborative learning and the other is intra-expert collaborative learning. The collaborative learning allows the knowledge transferring among different experts and image augmentations, which aims to reduce the learning uncertainties and promotes each expert model to achieve better performance. Our method is inherently a multi-expert framework. However, unlike previous works that need an ensemble of all experts to obtain the final predictions, our method can achieve the state-of-the-art performance only based on a single expert (better performance can be achieved based on the ensemble of all). This is because we collaboratively learn multiple experts by reducing the diversity of models rather than increasing their diversity as in previous works xiang2020learning ; wang2020long ; li2020overcoming ; cai2021ace ; zhang2021test . Our contributions can be summarized as follows:
@@ 2 Related Work
To alleviate the long-tailed class imbalance, lots of studies  zhou2020bbn ; cao2020domain ; xiang2020learning ; ren2020balanced ; zhang2021bag ; wang2020long have been conducted in recent years. The existing methods for long-tailed visual recognition can be roughly divided into three categories: class-rebalancing he2009learning ; buda2018systematic ; cui2019class ; huang2016learning ; ren2018learning ; wang2017learning , multi-stage training kang2019decoupling ; cao2019learning and multi-expert methods xiang2020learning ; wang2020long ; li2020overcoming ; cai2021ace ; zhang2021test . In the following, we will review the methods in the above three categories.
@@
Class re-balancing, which aims to re-balance the contribution of each class during training, is a classic and widely used method for long-tailed learning. Usually, there are two types for class re-balancing. The first one is data re-sampling chawla2002smote ; kang2019decoupling ; wang2020devil ; zhang2021learning ; zhang2021bag , including over-sampling chawla2002smote , under-sampling, square-root sampling kang2019decoupling and progressively-balanced sampling kang2019decoupling . The goal of data re-sampling is to increase the sampling probability of samples in tail classes while weakening that of samples in head classes. The second one is loss re-weighting lin2017focal ; wang2021seesaw ; ren2020balanced ; tan2020equalization ; zhao2018adaptive ; zhang2021bag , that is, re-weighting of loss function by the numbers of different classes. The popular methods in loss re-weighting include Focal loss lin2017focal , Seesaw loss wang2021seesaw , Balanced Softmax Cross-Entropy (BSCE) ren2020balanced , Class-Dependent Temperatures (CDT) ye2020identifying , LDAM loss cao2019learning and Equalization loss tan2020equalization . Class re-balancing improves the overall performance but usually at the sacrifice of the accuracy on head classes.
@@
Multi-stage training methods divide the training process into several stages hong2021disentangling ; menon2020long ; zhang2021distribution ; kang2019decoupling ; li2021self . For example, Kang et al. kang2019decoupling decouple the training procedure into representation learning and classifier learning, where the representation learning adopts the instance-balanced sampling to learn a good feature extractor and the classifier learning adopts the class-balanced sampling to re-adjust the classifier. Besides, some other works menon2020long ; zhang2021distribution ; hong2021disentangling ; zhao2021adaptive tend to improve performance via a post-process of shifting model logits. Li et al. li2021self propose a four-stage training strategy on basis of knowledge distillation, self-supervision and decoupled learning. To be specific, the network is first trained with classification and self-supervised losses, and then the feature extractor is frozen and the classifier is re-adjusted by the class-balanced sampling. Later, an additional network is employed to distill the knowledge from the previously trained network, and finally, the network is further trained by re-adjusting the classifier. As we can see, multi-stage training methods may rely on heuristic design.
@@
More recently, multi-expert frameworks xiang2020learning ; zhou2020bbn ; wang2020long ; zhang2021test ; cai2021ace ; li2020overcoming ; wang2020devil ; li2021trustworthy receive increasing concern, e.g., Learning From Multiple Expert (LFME) xiang2020learning , Bilateral-Branch Network (BBN) zhou2020bbn , RoutIng Diverse Experts (RIDE) wang2020long , Test-time Aggregating Diverse Experts (TADE) zhang2021test and Ally Complementary Experts (ACE) cai2021ace . For example, LFME xiang2020learning formulates three experts with each corresponding to one of head, medium and tail classes. RIDE wang2020long proposes a distribution-aware diversity loss to the multi-expert network and it encourages the experts to be diverse from each other. Multi-expert methods indeed improve the recognition accuracy for long-tailed learning, but those methods still need to be further exploited. For example, most current multi-expert methods employ different models to learn knowledge from different aspects, while the mutual supervision among them is deficient. Moreover, they often employ an ensemble to produce predictions, which leads to an increase in complexity.
@@
Knowledge distillation is a prevalent technology in knowledge transferring. One typical manner of knowledge distillation is teacher-student learning hinton2015distilling ; furlanello2018born , which transfers knowledge from a large teacher model to a small student model. Current methods in knowledge distillation can be divided into three categories: offline distillation hinton2015distilling ; romero2014fitnets ; passalis2018learning ; li2020few ; zagoruyko2016paying , online distillation zhang2018deep ; guo2020online ; chen2020online ; dvornik2019diversity ; chen2020online ; bhat2021distill and self-distillation zhang2020self ; zhang2019your ; hou2019learning ; phuong2019distillation ; yuan2019revisit . Early methods hinton2015distilling ; passalis2018learning often adopt an offline learning strategy, which transfers the knowledge from a pretrained teacher model to a student model. Most works hinton2015distilling ; passalis2018learning ; mirzadeh2020improved distill the knowledge from the output distributions, while some works achieve the knowledge transferring by matching feature representations romero2014fitnets or attention maps zagoruyko2016paying . The offline distillation is very popular in the early stage. However, the offline way only considers transferring the knowledge from the teacher to the student, and therefore, the teacher normally should be a more complex high-capacity model than the student. In recent years, knowledge distillation has been extended to an online way zhang2018deep ; guo2020online ; chen2020online ; dvornik2019diversity ; chen2020online ; walawalkar2020online , where the whole knowledge distillation is conducted in an one-phase and end-to-end training scheme. For example, in Deep Mutual Learning zhang2018deep , any one model can be a student and can distill knowledge from all other models. Zhu et al. lan2018knowledge propose a multi-branch architecture with treating each branch as a student to further reduce computational cost. Compared with offline distillation, online distillation is more efficient by taking the learning in an one-phase end-to-end scheme. For self-distillation zhang2020self ; zhang2019your ; hou2019learning ; phuong2019distillation ; yuan2019revisit , it can be regarded as a special case in online distillation, where the teacher and the student refer to the same network. In other words, self-distillation means that the model distills the knowledge from itself. For example, Zhang et al. zhang2019your divide the network into several sections according to their depth, and allow the low sections to distill the knowledge from high sections. Our work aims to reduce the predictive uncertainty in long-tailed learning by online distillation and self-distillation. However, different from previous works, we formulate the distillation in a nested way to focus on all categories as well as some important categories, which facilitates the network to obtain the meticulous distinguishing ability.
@@ 3.1 Balanced Individual Learning
where $z_{ij}^{k}$ is the model’s class-$j$ output and $C$ is the number of classes. This is a widely used way to compute the predicted probability, and some losses like Cross Entropy (CE) loss is computed based on it. However, this way does not consider the data distribution, and is not suitable for long-tailed visual recognition, where a vanilla model based on ${\tilde{\bf p}}({\bf x}_{i};{\bm{\theta}}_{k})$ would be largely dominated by head classes. Therefore, some researchers ren2020balanced proposed to compute predicted probability of class-$j$ in a balanced way:
@@
where $n_{j}$ is the total number of samples of class $j$. In this way, contributions of tail classes are strengthened while contributions of head classes are suppressed. Based on such balanced probabilities, Ren et al. ren2020balanced further proposed a Balanced Softmax Cross Entropy (BSCE) loss to alleviate long-tailed class imbalance in model training. We also adopt the BSCE to conduct the individual learning for each expert, which ensures that each one can achieve the strong discrimination ability. Mathematically, the loss of the balanced individual learning for expert-$k$ can be denoted as:

### v3

@@ Abstract
Long-tailed visual recognition has received increasing attention in recent years. Due to the extremely imbalanced data distribution in long-tailed learning, the learning process shows great uncertainties. For example, the predictions of different experts on the same image vary remarkably despite the same training settings. To alleviate the uncertainty, we propose a Nested Collaborative Learning (NCL++) which tackles the long-tailed learning problem by a collaborative learning. To be specific, the collaborative learning consists of two folds, namely inter-expert collaborative learning (InterCL) and intra-expert collaborative learning (IntraCL). InterCL learns multiple experts collaboratively and concurrently, aiming to transfer the knowledge among different experts. IntraCL is similar to InterCL, but it aims to conduct the collaborative learning on multiple augmented copies of the same image within the single expert. To achieve the collaborative learning in long-tailed learning, the balanced online distillation is proposed to force the consistent predictions among different experts and augmented copies, which reduces the learning uncertainties. Moreover, in order to improve the meticulous distinguishing ability on the confusing categories, we further propose a Hard Category Mining (HCM), which selects the negative categories with high predicted scores as the hard categories. Then, the collaborative learning is formulated in a nested way, in which the learning is conducted on not just all categories from a full perspective but some hard categories from a partial perspective. Extensive experiments manifest the superiority of our method with outperforming the state-of-the-art whether with using a single model or an ensemble. The code is available at https://github.com/Bazinga699/NCL.
@@ 1 Introduction
In recent years, deep neural networks have achieved resounding success in various visual tasks, e.g., image classification [1, 2], object detection [3, 4], semantic segmentation [5, 6] and so on. Despite the advances in deep technologies and computing capability, the huge success also highly depends on large well-designed datasets of having a roughly balanced distribution, such as ImageNet [7], MS COCO [8] and Places [9]. This differs notably from real-world datasets, which usually exhibits long-tailed data distributions [10, 11] where few head classes occupy most of the data while many tail classes only have few samples as shown in Fig. 1. In such scenarios, the model is easily dominated by those few head classes, whereas low accuracy rates are usually achieved for many other tail classes. Undoubtedly, the long-tailed characteristics challenge deep visual recognition, and also immensely hinder the practical use of deep models.
In long-tailed visual recognition, most early works focus on designing the class re-balancing strategies [12, 13, 14, 15, 16, 10] and decoupled learning [17, 18]. Those methods can accomplish some accuracy improvements but still cannot deal with the long-tailed class imbalance problem well. For example, class re-balancing methods are often confronted with risk of overfitting. More recent efforts aim to improve the long-tailed learning by using multiple experts [19, 20, 21, 22, 23]. The multi-expert algorithms follow a straightforward idea of making them diverse from each other. To achieve this, some works [19, 23] force different experts to focus on different aspects. For example, LFME [19] formulates a network with three experts and it forces each expert to learn samples from one of head, middle and tail classes. Besides, there are some works [20] adopt some constraint losses to diversify the multiple experts. For example, RIDE [20] proposes a distribution-aware diversity loss to achieve this. The previous multi-expert methods can achieve good performance mainly due to the diversity among them. Also because of this, these methods obtain the predictions through the ensemble rather than a single expert due to that only employing a single expert hardly achieves reliable recognition.
@@
However, we found that there are great uncertainties in the long-tailed learning. The learning uncertainties mainly come from two aspects. One is that the same expert performs differently for the same image with different augmentations. The other is that, for two experts with the same configuration and the same training settings, they still vary greatly on predictions with respect to the same input image. To analyze this, we visualize the Kullback-Leibler (KL) distance [^1] https://en.wikipedia.org/wiki/Kullback-Leibler_divergence over the predictions generated by two experts with the same training settings, and the predictions generated by two augmented image copies with respect to the same expert. As shown in Fig. 2 (b), the green color shows the KL distance distribution for two different experts of the same structure and training settings with taking the same image as the input, and the blue color shows that of the same expert with taking different augmentations of the same image. Specifically, to analyze the uncertainties across all categories, the Kullback-Leibler (KL) distance is specifically calculated for all categories with reporting the average distance over all images for each class. As we can see, the predictions vary greatly especially in tail classes, which signifies the great uncertainty in the learning process as well as the diversity in models. To show the prediction differences in details, we visualize the detailed predictions of the two experts with the same structure and training settings on a randomly selected sample (corresponding to the green color in Fig. 2 (b)). According to previous works [19, 20, 21, 22, 23], a simple ensemble like averaging the predictions could be employed to reduce the uncertainty in the prediction stage, which can achieve a certain performance improvement. However, an ensemble of multiple experts also brings huge amount of calculation and parameters, which limits its practical use. This brings up a question, can we utilize these uncertainties to improve the performance of single model?
@@
Previous works [24, 25, 26] simply conduct the classification and distillation from a full perspective on all categories. It helps the network to obtain the global discrimination on all categories, but lacks of meticulous distinguishing ability on the confusing categories. In the classification problem, the network should pay attention to the hard negative categories with high predictive scores rather than simply treating all categories equally, which helps to reduce the confusion between the target category and the confusing categories. To achieve this, we first propose a Hard Category Mining (HCM) to select the hard categories for each sample, in which the hard category is defined as the negative category with a high predictive score. Then, we formulate a nested learning for the supervised individual learning of a single network as well as the collaborative learning (including IntraCL and InterCL). To the specific, the nested learning conducts the supervised learning or distillation from two perspectives, where one is the full perspective on all categories and the other is the partial perspective only on the selected hard categories. In this way, not only the global discriminative capability but also the meticulous distinguishing ability can be captured.
The proposed method utilizes the collaborative learning for long-tailed visual recognition. The proposed collaborative learning is two folds, where one is inter-expert collaborative learning and the other is intra-expert collaborative learning. The collaborative learning allows the knowledge transferring among different experts and image augmentations, which aims to reduce the learning uncertainties and promotes each expert model to achieve better performance. Our method is inherently a multi-expert framework. However, unlike previous works that need an ensemble of all experts to obtain the final predictions, our method can achieve the state-of-the-art performance only based on a single expert (better performance can be achieved based on the ensemble of all). This is because we collaboratively learn multiple experts by reducing the diversity of models rather than increasing their diversity as in previous works [19, 20, 21, 22, 23]. Our contributions can be summarized as follows:
@@ 2 Related Work
To alleviate the long-tailed class imbalance, lots of studies  [27, 28, 19, 29, 30, 20] have been conducted in recent years. The existing methods for long-tailed visual recognition can be roughly divided into three categories: class-rebalancing [12, 13, 14, 15, 16, 10], multi-stage training [17, 18] and multi-expert methods [19, 20, 21, 22, 23]. In the following, we will review the methods in the above three categories.
@@
Class re-balancing, which aims to re-balance the contribution of each class during training, is a classic and widely used method for long-tailed learning. Usually, there are two types for class re-balancing. The first one is data re-sampling [31, 17, 32, 33, 30], including over-sampling [31], under-sampling, square-root sampling [17] and progressively-balanced sampling [17]. The goal of data re-sampling is to increase the sampling probability of samples in tail classes while weakening that of samples in head classes. The second one is loss re-weighting [34, 35, 29, 36, 37, 30], that is, re-weighting of loss function by the numbers of different classes. The popular methods in loss re-weighting include Focal loss [34], Seesaw loss [35], Balanced Softmax Cross-Entropy (BSCE) [29], Class-Dependent Temperatures (CDT) [38], LDAM loss [18] and Equalization loss [36]. Class re-balancing improves the overall performance but usually at the sacrifice of the accuracy on head classes.
@@
Multi-stage training methods divide the training process into several stages [39, 40, 41, 17, 42]. For example, Kang et al. [17] decouple the training procedure into representation learning and classifier learning, where the representation learning adopts the instance-balanced sampling to learn a good feature extractor and the classifier learning adopts the class-balanced sampling to re-adjust the classifier. Besides, some other works [40, 41, 39, 43] tend to improve performance via a post-process of shifting model logits. Li et al. [42] propose a four-stage training strategy on basis of knowledge distillation, self-supervision and decoupled learning. To be specific, the network is first trained with classification and self-supervised losses, and then the feature extractor is frozen and the classifier is re-adjusted by the class-balanced sampling. Later, an additional network is employed to distill the knowledge from the previously trained network, and finally, the network is further trained by re-adjusting the classifier. As we can see, multi-stage training methods may rely on heuristic design.
@@
More recently, multi-expert frameworks [19, 27, 20, 23, 22, 21, 32, 44] receive increasing concern, e.g., Learning From Multiple Expert (LFME) [19], Bilateral-Branch Network (BBN) [27], RoutIng Diverse Experts (RIDE) [20], Test-time Aggregating Diverse Experts (TADE) [23] and Ally Complementary Experts (ACE) [22]. For example, LFME [19] formulates three experts with each corresponding to one of head, medium and tail classes. RIDE [20] proposes a distribution-aware diversity loss to the multi-expert network and it encourages the experts to be diverse from each other. Multi-expert methods indeed improve the recognition accuracy for long-tailed learning, but those methods still need to be further exploited. For example, most current multi-expert methods employ different models to learn knowledge from different aspects, while the mutual supervision among them is deficient. Moreover, they often employ an ensemble to produce predictions, which leads to an increase in complexity.
@@
Knowledge distillation is a prevalent technology in knowledge transferring. One typical manner of knowledge distillation is teacher-student learning [45, 46], which transfers knowledge from a large teacher model to a small student model. Current methods in knowledge distillation can be divided into three categories: offline distillation [45, 47, 48, 49, 50], online distillation [26, 24, 51, 52, 51, 53] and self-distillation [54, 55, 56, 57, 58]. Early methods [45, 48] often adopt an offline learning strategy, which transfers the knowledge from a pretrained teacher model to a student model. Most works [45, 48, 59] distill the knowledge from the output distributions, while some works achieve the knowledge transferring by matching feature representations [47] or attention maps [50]. The offline distillation is very popular in the early stage. However, the offline way only considers transferring the knowledge from the teacher to the student, and therefore, the teacher normally should be a more complex high-capacity model than the student. In recent years, knowledge distillation has been extended to an online way [26, 24, 51, 52, 51, 60], where the whole knowledge distillation is conducted in an one-phase and end-to-end training scheme. For example, in Deep Mutual Learning [26], any one model can be a student and can distill knowledge from all other models. Zhu et al. [25] propose a multi-branch architecture with treating each branch as a student to further reduce computational cost. Compared with offline distillation, online distillation is more efficient by taking the learning in an one-phase end-to-end scheme. For self-distillation [54, 55, 56, 57, 58], it can be regarded as a special case in online distillation, where the teacher and the student refer to the same network. In other words, self-distillation means that the model distills the knowledge from itself. For example, Zhang et al. [55] divide the network into several sections according to their depth, and allow the low sections to distill the knowledge from high sections. Our work aims to reduce the predictive uncertainty in long-tailed learning by online distillation and self-distillation. However, different from previous works, we formulate the distillation in a nested way to focus on all categories as well as some important categories, which facilitates the network to obtain the meticulous distinguishing ability.
@@ 3.1 Balanced Individual Learning
where $z_{ij}^{k}$ is the model’s class-$j$ output and $C$ is the number of classes. This is a widely used way to compute the predicted probability, and some losses like Cross Entropy (CE) loss is computed based on it. However, this way does not consider the data distribution, and is not suitable for long-tailed visual recognition, where a vanilla model based on ${\tilde{\bf p}}({\bf x}_{i};{\bm{\theta}}_{k})$ would be largely dominated by head classes. Therefore, some researchers [29] proposed to compute predicted probability of class-$j$ in a balanced way:
@@
where $n_{j}$ is the total number of samples of class $j$. In this way, contributions of tail classes are strengthened while contributions of head classes are suppressed. Based on such balanced probabilities, Ren et al. [29] further proposed a Balanced Softmax Cross Entropy (BSCE) loss to alleviate long-tailed class imbalance in model training. We also adopt the BSCE to conduct the individual learning for each expert, which ensures that each one can achieve the strong discrimination ability. Mathematically, the loss of the balanced individual learning for expert-$k$ can be denoted as: